Jerlyn and Mico have a stamp. Jerlyn has 1 2/3 of the number of Mico's stamps. If they have 120 stamps altogether, find the number of stamps that Mico has.

Answer by stanbon(75887)

Jerlyn has 1 2/3 of the number of Mico's stamps. If they have 120 stamps altogether, find the number of stamps that Mico has.
-------
J = (5/3)M
J + M = 120
--------------------
Substitute for "J" and solve for "M":
(5/3)M + M = 120
(8/3)M = 120
----
M = (3/8)120
M = 3*15 = 45 (# of stamps Mico has)
---------------
Cheers,
Stan H.
----------